# Secure and Efficient Access Control Framework for Computer-Use Agents via Context Space

### v2

@@ 4.3. LLM-based Context Analyzer
Challenges in GUI analysis. For API- and CLI-based apps, policy generation is straightforward due to the availability of well-defined function specifications. Specifically, API tools usually provide detailed documentation on purpose, parameters, and behavior (Qin et al., 2023; LangChain, 2025c), while CLIs usually offer manuals specifying command functionality and usage (Kerrisk, 2025; Microsoft, 2025b), which enables LLMs to generate high-quality context spaces with sufficient semantic information on function behavior. However, GUI agents present greater challenges. Concretely, GUI agents interact with apps through generic interaction primitives (such as screen capture, clicking, scrolling, and text input) rather than explicit functional interfaces. Even worse, GUI apps typically lack functional documentation. While human users understand app usage through interaction, this knowledge is rarely formalized in machine-readable specifications, making it difficult to identify the security-sensitive functions that require protection. As a result, no existing agent protection approach has demonstrated automated end-to-end policy generation for GUI apps. Existing approaches for GUI functionality discovery include static program analysis and GUI exploration tools. However, our experiments (§7.2) reveal limitations in both approaches: static analysis often fails to capture the complete scope of GUI functionalities due to dynamic binding and complex UI frameworks, while GUI exploration tools suffer from scalability issues and frequently encounter infinite loops or state explosion problems that prevent comprehensive coverage.
@@ 4.4. Optimized Context Manager
Challenges in performance. The context manager faces two primary performance challenges. First, LLM-based context extraction from user instructions introduces additional inference latency. One simple approach is to request the LLM to analyze the relevant intent and extract context values after the agent determines the function to perform, which would introduce inference latency for every agent action, severely degrading performance (Shi et al., 2025a; Lee et al., 2025). Second, context management for complex apps creates substantial computational overhead. Large-scale apps, particularly GUI apps, often have extensive context spaces involving numerous contexts. Frequent acquisition and updates of all context values incur unnecessary computational costs, increased resource consumption, and additional energy drain. Moreover, loading these context spaces is time-consuming, especially when switching between multiple apps, resulting in a poor user experience.
@@ 5. Implementation
Automatic policy generation. As outlined in §4.3, we use LLMs (DeepSeek-R1 (Guo et al., 2025)) for policy generation. For API and CLI agents, the LLM analyzes function documentation to generate policies. For GUI agents, we target Android apps since Android is widely used and open-source. For a given app’s source code, our context analyzer first uses text matching to identify files related to GUI operations, then provides these files to the LLM for analysis of event handlers. The LLM outputs results in JSON format, including GUI element details like class and resource ID, and methods invoked by the corresponding handlers. We then use static analysis tools (CodeQL (GitHub, 2021)) to construct call graphs.
@@
Context acquisition and policy validation. Context values come from user instructions, device systems, and agent operations. During intent selection, the intent extractor LLM parses user requests and environmental data (e.g., configurations) to obtain contexts. Agent-related contexts are captured in two ways: parameter contexts are obtained when intercepting agent actions, while history contexts are updated after action completion. For system-level contexts, we currently support CLI-based acquisition via tools like ADB (Google, 2025b). Another approach is implementing dedicated system services to provide unified interfaces for context acquisition, which requires platform vendor participation in the ecosystem.
@@ 7.1. Experimental Setup and Benchmarks (RQ1)
To evaluate CSAgent across different agent interaction modalities, we employ three benchmarks: AgentBench  (Liu et al., 2023) (OS environment), AgentDojo  (Debenedetti et al., 2024), and AndroidWorld  (Rawles et al., 2024), which respectively evaluate CLI-based, API-based, and GUI-based agents. We employ DeepSeek-V3  (Liu et al., 2024b) as our runtime text model, and Seed1.6  (ByteDance, 2025) as the multimodal model. The integration process with these benchmarks closely mirrors real-world CSAgent deployment, consisting of two steps. First, we analyze the applications within each benchmark, including shell commands, tools, and apps, to generate context spaces using our context analyzer. Second, we adapt the agent frameworks to the CSAgent service with minimal code modifications: establishing an RPC connection, invoking the intent extraction method upon user request arrival, and calling validation and context update methods before and after agent actions. Our successful support of these benchmarks demonstrates CSAgent’s strong compatibility (satisfying G4).
Our evaluation methodology encompasses three primary dimensions across all benchmarks: utility, security, and overhead. Utility measures the agent’s task completion capability under CSAgent protection, security evaluates CSAgent’s ability to block unintended behaviors, and overhead quantifies the latency and token consumption introduced by CSAgent. For AndroidWorld specifically, we also conduct an analysis of CSAgent’s code identification capabilities for context space generation. We compare it against UI-CTX  (Li et al., 2025a) and AutoDroid  (Wen et al., 2024), representing state-of-the-art approaches in static analysis and GUI exploration, respectively. For security evaluation, all the context policies are generated by the context analyzer (satisfying G3). While AgentDojo provides built-in adversarial tests, we construct additional security test cases for AgentBench and AndroidWorld. These test cases consist of mismatched user requests and agent actions (i.e., simulating agent actions that violate the user intent), creating a series of pairs that serve as input to evaluate capability to detect anomalous agent behaviors that deviate from user intentions. Notably, the policies we test are generated solely based on the apps themselves, and the policy generation process has no access to the test set.
To emphasize effectiveness, we design our experiments around CSAgent and two other agent configurations. First, we evaluate a Vanilla Agent that operates without any security protection mechanisms, serving as our analysis baseline. Second, we implement a Pre-execution Validation Agent (PVAgent) that employs dynamic policy generation and validation before executing each action (similar to (Tsai and Bagdasarian, 2025; Shi et al., 2025a)).
@@ 7.2. Capability of GUI Analysis (RQ2)
Figure 6 demonstrates CSAgent’s capability in extracting GUI elements and event handlers from open-source apps and AOSP system apps used by AndroidWorld. UI-CTX was configured with default settings, while AutoDroid employed the deep learning-based strategy Humanoid (Li et al., 2019). The results show that CSAgent extracts 0.93× more GUI element information (including handlers) than AutoDroid and 3.12× more than UI-CTX (satisfying G3). While some GUI elements are identified by these two baseline methods but missed by CSAgent, most do not produce actual control effects (e.g., merely opening menus rather than clicking specific options). We prompt our context analyzer to ignore these non-functional elements to optimize token consumption. We did not compare to source code-based analysis primarily due to the lack of available tools. Our approach differs by using LLMs to identify GUI-related code, offering better generalization capabilities. Importantly, CSAgent can complement other methods to achieve more comprehensive coverage.
@@ 7.3. Effectiveness of CSAgent Access Control (RQ3)
Agent
Type
AgentDojo AgentBench AndroidWorld
@@
Vanilla 47.18% 53.75% - -
PVAgent 4.82% 33.21% 1.00% 5.09%
CSAgent 1.97% 35.48% 0.11% 1.21%
@@
Agent
Type
Banking Slack Travel Workspace
@@
Vanilla 45.14 59.72 85.71 64.76 51.43 27.14 6.43 63.39
PVAgent 0.00 36.81 0.00 4.76 18.57 40.71 0.71 50.54
CSAgent 0.00 40.97 0.00 18.10 7.86 32.14 0.00 50.71
Table 3 presents CSAgent’s comprehensive attack defense capabilities. The primary evaluation metric is Attack Success Rate (ASR), which measures the success rate of simulated attacks against agents. Additionally, AgentDojo includes the Utility Under Attacks (UUA) metric, which evaluates an agent’s ability to complete original tasks under injection attacks. The results demonstrate that CSAgent achieves a geometric mean ASR of only 0.64%, successfully defending against approximately 99.36% of attacks (satisfying G1). In comparison, PVAgent defends against 97.09% of attacks. Table 4 provides detailed results for AgentDojo’s prompt injection attack tests. Except for the travel test set, CSAgent achieves 100% successful defense across all scenarios. Our analysis reveals that these successful attacks are content safety-related, which is beyond the scope. For AgentBench and AndroidWorld, CSAgent did not defend against all attacks. This is due to the context analyzer generating policies that were not sufficiently comprehensive in a single analysis. After a round of iteration through the PEF, CSAgent was able to defend against all attacks. CSAgent outperforms PVAgent mainly due to the nature of static policies, allowing us to use stronger reasoning models to generate and iteratively improve policies with low impact on runtime performance.
@@ 7.4. Performance and Cost Overhead (RQ4)
Figure 7 illustrates the utility drop and extra latency introduced by CSAgent. On AgentBench and AndroidWorld, both utility drop and latency overhead introduced by CSAgent are less than 5%, compared to PVAgent’s overhead of 58.57% and 34.28%. On AgentDojo, CSAgent introduces 48.46% performance overhead, compared to PVAgent’s 103.98% overhead, while a more complex dynamic policy-based method incurs a 1.94× slowdown (Shi et al., 2025a). Additionally, we found that 80.08% of contexts are classified as cold or warm temperature levels. These results demonstrate that our optimized context manager efficiently performs user intent extraction, context updates and policy verification (satisfying G2). For utility, CSAgent causes a 9.33% decrease on average. The degradation on AgentDojo is more pronounced, primarily due to task execution dependencies on untrusted information sources, as discussed earlier. Notably, we terminate a task once policy validation fails during evaluation. In real deployments, user intervention or agent reflection could resume such tasks, further reducing the impact. We also evaluated CSAgent’s impact on task completion steps. Overall, the impact is minimal, indicating that CSAgent’s decisions to allow or block agent actions are relatively accurate.
@@
Table 5 presents CSAgent’s per task token consumption compared to the vanilla agent. The geometric mean of consumption overhead is 35.30%, with over 96.41% consisting of prompt tokens, which are less expensive than completion tokens. Generally, more complex tasks have proportionally lower overhead. For instance, AgentBench tasks have low average token consumption, so the large context space results in higher overhead. Conversely, AndroidWorld tasks are highly complex, requiring multimodal capabilities for screen recognition, leading to very high task token consumption and lower overhead. In comparison, runtime policy generation methods introduce 1.61× additional token consumption on AgentDojo (Shi et al., 2025a). CSAgent’s low token consumption results from development-phase policy generation that applies across all devices and users without regeneration, and from requiring only one additional LLM inference per task regardless of complexity (satisfying G2).
@@ 9. Related Work
Access control and contextual integrity. Traditional access control mechanisms in OS include discretionary access control (DAC) and mandatory access control (MAC) (Sandhu and Samarati, 1996). SELinux  (Smalley and Craig, 2013; Wright et al., 2002) exemplifies MAC principles by defining fine-grained policies that govern process interactions and resource access based on security contexts. The theory of contextual integrity (CI) (Nissenbaum, 2004, 2009), proposed in more recent years, has been explored by researchers for its applications in existing operating systems (Roesner et al., 2012; Wijesekera et al., 2015), driving the evolution of permission management and access control toward more user-transparent, secure, and flexible approaches.
@@
Access control for LLM agents. For LLM agents, access control focuses on constraining tool calls through pre-execution validation. Most approaches use contextual information to make authorization decisions, following the CI principle. AirGapAgent (Bagdasarian et al., 2024) employs a separate LLM to directly make decisions about data access permissions. GuardAgent (Xiang et al., 2024) employs LLMs to generate guardrail code that validates predefined rules before executing actions. VeriSafe Agent (Lee et al., 2025) describes rules through Horn clauses and verifies agent behavior based on logic-based reasoning. Conseca (Tsai and Bagdasarian, 2025) and Progent (Shi et al., 2025a) achieve dynamic security policy generation, enabling more autonomous responses to diverse task scenarios. AgentSentinel (Hu et al., 2025) extends context extraction to the OS, providing real-time protection for CUAs by combining rule-based and LLM-based auditing. These approaches rely on runtime LLM-based security decision-making, suffering from security or performance limitations. AgentSpec (Wang et al., 2025b) employs predefined static policies, but adopts a risk-blocking paradigm that struggles with unknown threats and lacks user intent differentiation, limiting policy flexibility.
@@
Other system-level LLM agent protection. Beyond access control, some other forms of system-level protection have been proposed for agents. The Instruction Hierarchy  (Wallace et al., 2024) and StruQ  (Chen et al., 2024) implement privilege and type separation for LLM inputs to mitigate interference between different input sources. f-secure LLM system  (Wu et al., 2024), RTBAS  (Zhong et al., 2025), and SAFEFLOW  (Li et al., 2025c) introduce information flow control into LLM agents to prevent prompt injection and privacy leakage. GoEX  (Patil et al., 2024b) implements post-facto validation of agent actions and isolates dangerous operations through sandboxing. IsolateGPT  (Wu et al., 2025) employs isolation for multi-LLM application scenarios to prevent interference between LLMs and different applications. ACE (Li et al., 2025b) decouples agent planning into two phases to defend against untrusted third-party apps. These approaches and CSAgent can collectively form part of a defense-in-depth strategy (Google, 2025c), working in conjunction to achieve more comprehensive protection.

### v3

@@ 4.3. LLM-based Context Analyzer
Challenges in GUI analysis. For API- and CLI-based apps, policy generation is straightforward due to the availability of well-defined function specifications. Specifically, API tools usually provide detailed documentation on purpose, parameters, and behavior (qin2023toolllm; website:langchaintool), while CLIs usually offer manuals specifying command functionality and usage (website:linuxman; website:powershelldoc), which enables LLMs to generate high-quality context spaces with sufficient semantic information on function behavior. However, GUI agents present greater challenges. Concretely, GUI agents interact with apps through generic interaction primitives (such as screen capture, clicking, scrolling, and text input) rather than explicit functional interfaces. Even worse, GUI apps typically lack functional documentation. While human users understand app usage through interaction, this knowledge is rarely formalized in machine-readable specifications, making it difficult to identify the security-sensitive functions that require protection. As a result, no existing agent protection approach has demonstrated automated end-to-end policy generation for GUI apps. Existing approaches for GUI functionality discovery include static program analysis and GUI exploration tools. However, our experiments (§7.2) reveal limitations in both approaches: static analysis often fails to capture the complete scope of GUI functionalities due to dynamic binding and complex UI frameworks, while GUI exploration tools suffer from scalability issues and frequently encounter infinite loops or state explosion problems that prevent comprehensive coverage.
@@ 4.4. Optimized Context Manager
Challenges in performance. The context manager faces two primary performance challenges. First, LLM-based context extraction from user instructions introduces additional inference latency. One simple approach is to request the LLM to analyze the relevant intent and extract context values after the agent determines the function to perform, which would introduce inference latency for every agent action, severely degrading performance (shi2025progent; lee2025verisafe). Second, context management for complex apps creates substantial computational overhead. Large-scale apps, particularly GUI apps, often have extensive context spaces involving numerous contexts. Frequent acquisition and updates of all context values incur unnecessary computational costs, increased resource consumption, and additional energy drain. Moreover, loading these context spaces is time-consuming, especially when switching between multiple apps, resulting in a poor user experience.
@@ 5. Implementation
Automatic policy generation and evolution. As outlined in §4.3, we use LLMs (DeepSeek-R1 (guo2025deepseek)) for policy generation and refinement. For API and CLI agents, the LLM analyzes function documentation to generate policies. For GUI agents, we target Android apps since Android is widely used and open-source. For a given app’s source code, our context analyzer first uses text matching to identify files related to GUI operations, then provides these files to the LLM for analysis of event handlers. The LLM outputs results in JSON format, including GUI element details like class and resource ID, and methods invoked by the corresponding handlers. We then use static analysis tools (CodeQL (website:codeql)) to construct call graphs. For policy evolution, we implement a runtime logging system that captures task information, validation events, user feedback, and execution traces. During refinement, the LLM analyzes the logs alongside the existing context space and update history to identify potential utility and security issues, then systematically updates policies to address these deficiencies.
@@
Context acquisition and policy validation. Context values come from user instructions, device systems, and agent operations. During intent selection, the intent extractor LLM parses user requests and environmental data (e.g., configurations) to obtain contexts. Agent-related contexts are captured in two ways: parameter contexts are obtained when intercepting agent actions, while history contexts are updated after action completion. For system-level contexts, we currently support CLI-based acquisition via tools like ADB (website:adb). Another approach is implementing dedicated system services to provide unified interfaces for context acquisition, which requires platform vendor participation in the ecosystem.
@@ 7. Evaluation
RQ5: How effectively does the Policy Evolution Framework refine context spaces to optimize the trade-off between security, utility, and performance?
@@ 7.1. Experimental Setup and Benchmarks (RQ1)
To evaluate CSAgent across different agent interaction modalities, we employ three benchmarks: AgentBench  (liu2023agentbenchevaluatingllmsagents) (OS environment), AgentDojo  (agentdojo), and AndroidWorld  (rawles2024androidworld), which respectively evaluate CLI-based, API-based, and GUI-based agents. We employ DeepSeek-V3  (liu2024deepseekv3) as our runtime text model, and Seed1.6  (website:seed16) as the multimodal model. The integration process with these benchmarks closely mirrors real-world CSAgent deployment, consisting of two steps. First, we analyze the applications within each benchmark, including shell commands, tools, and apps, to generate context spaces using our context analyzer. Second, we adapt the agent frameworks to the CSAgent service with minimal code modifications: establishing an RPC connection, invoking the intent extraction method upon user request arrival, and calling validation and context update methods before and after agent actions. Our successful support of these benchmarks demonstrates CSAgent’s strong compatibility (satisfying G4).
Our evaluation methodology encompasses three primary dimensions across all benchmarks: utility, security, and overhead. Utility measures the agent’s task completion capability under CSAgent protection, security evaluates CSAgent’s ability to block unintended behaviors, and overhead quantifies the latency and token consumption introduced by CSAgent. For AndroidWorld specifically, we also conduct an analysis of CSAgent’s code identification capabilities for context space generation. We compare it against UI-CTX  (li2025uictx) and AutoDroid  (wen2024autodroid), representing state-of-the-art approaches in static analysis and GUI exploration, respectively. For security evaluation, all the context policies are generated by the context analyzer (satisfying G3). While AgentDojo provides built-in adversarial tests, we construct additional security test cases for AgentBench and AndroidWorld. These test cases consist of mismatched user requests and agent actions (i.e., simulating agent actions that violate the user intent), creating a series of pairs that serve as input to evaluate capability to detect anomalous agent behaviors that deviate from user intentions. Notably, the policies we test are generated solely based on the apps themselves, and the policy generation process has no access to the test set.
To emphasize effectiveness, we design our experiments around CSAgent and two other agent configurations. First, we evaluate a Vanilla Agent that operates without any security protection mechanisms, serving as our analysis baseline. Second, we implement a Pre-execution Validation Agent (PVAgent) that employs dynamic policy generation and validation before executing each action (similar to (tsai2025contextualagent; shi2025progent)). Finally, we evaluate CSAgent-RF, a refined variant of CSAgent that leverages PEF, to further demonstrate its effectiveness.
@@ 7.2. Capability of GUI Analysis (RQ2)
Figure 6 demonstrates CSAgent’s capability in extracting GUI elements and event handlers from open-source apps and AOSP system apps used by AndroidWorld. UI-CTX was configured with default settings, while AutoDroid employed the deep learning-based strategy Humanoid (li2020humanoiddeeplearningbasedapproach). The results show that CSAgent extracts 0.93× more GUI element information (including handlers) than AutoDroid and 3.12× more than UI-CTX (satisfying G3). While some GUI elements are identified by these two baseline methods but missed by CSAgent, most do not produce actual control effects (e.g., merely opening menus rather than clicking specific options). We prompt our context analyzer to ignore these non-functional elements to optimize token consumption. We did not compare to source code-based analysis primarily due to the lack of available tools. Our approach differs by using LLMs to identify GUI-related code, offering better generalization capabilities. Importantly, CSAgent can complement other methods to achieve more comprehensive coverage.
@@ 7.3. Effectiveness of CSAgent Access Control (RQ3)
Agent Type AgentDojo AgentBench AndroidWorld
@@
Vanilla 46.61% 54.26% - -
PVAgent 4.14% 33.03% 1.00% 5.09%
CSAgent 0% 41.61% 0.11% 1.21%
CSAgent-RF 0% 54.23% 0% 0%
@@
Agent Type Banking Slack Travel Workspace
@@
Vanilla 45.14 59.72 85.71 64.76 49.17 29.17 6.43 63.39
PVAgent 0.00 36.81 0.00 4.76 15.83 40.00 0.71 50.54
CSAgent 0.00 40.97 0.00 18.10 0.00 40.83 0.00 50.71
CSAgent-RF 0.00 56.25 0.00 38.10 0.00 54.17 0.00 68.39
Table 3 presents CSAgent’s comprehensive attack defense capabilities. The primary evaluation metric is Attack Success Rate (ASR), which measures the success rate of simulated attacks against agents. Additionally, AgentDojo includes the Utility Under Attacks (UUA) metric, which evaluates an agent’s ability to complete original tasks under injection attacks. For the AgentDojo travel task, we exclude one out-of-scope attack that targets LLM output content rather than agent behavior. The results demonstrate that CSAgent achieves average ASR of only 0.44%, successfully defending against approximately 99.56% of attacks (satisfying G1). In comparison, PVAgent defends against 96.59% of attacks. Table 4 provides detailed results for AgentDojo’s prompt injection attack tests. CSAgent achieves 100% successful defense across all scenarios. For AgentBench and AndroidWorld, CSAgent did not defend against all attacks. This is due to the context analyzer generating policies that were not sufficiently comprehensive in a single analysis. After a round of iteration through the PEF, CSAgent was able to defend against all attacks. CSAgent outperforms PVAgent mainly due to the nature of static policies, allowing us to use stronger reasoning models to generate and iteratively improve policies with low impact on runtime performance.
@@
The observed UUA drop is due to the high complexity of AgentDojo tasks and the information gap during initial policy generation. Specifically, many tasks in AgentDojo are derived from environmental contexts (e.g., billing files or to-do lists) or depend on implicit intermediate actions, rather than direct user instructions. Furthermore, some tasks rely on untrusted data sources, from which CSAgent restricts context extraction to ensure security. In our setup, local context sources (e.g., file systems) are configured as trusted, while online sources (e.g., web and email) are treated as untrusted. Since our context analyzer generates policies based solely on tool code without access to specific task scenarios, the LLM cannot comprehensively reason about such complex execution patterns or diverse user intents in a single analysis, leading to runtime failure. However, as shown in §7.5, the PEF effectively addresses this limitation through iterative refinement based on runtime feedback.
@@ 7.4. Performance and Cost Overhead (RQ4)
Figure 7 illustrates the utility drop and extra latency introduced by CSAgent. On all three benchmarks, both utility drop and latency overhead introduced by CSAgent are less than 5%, compared to PVAgent’s overhead of 58.57%, 34.28%, and 103.98% on AgentBench, AndroidWorld, and AgentDojo, respectively. Additionally, we found that 80.08% of contexts are classified as cold or warm temperature levels. These results demonstrate that our optimized context manager efficiently performs user intent extraction, context updates and policy verification (satisfying G2). For utility, CSAgent causes a 15.58% decrease on average. The degradation on AgentDojo is more pronounced, primarily due to task complexity and insufficient policy quality, as discussed earlier. Notably, upon a policy validation failure during evaluation, the agent is prompted to either explore alternative execution paths or terminate the task. In real deployments, user intervention or agent reflection could resume such tasks, further reducing the impact. To explore this potential, we evaluate a variant incorporating agent reflection upon validation failures. On AgentDojo, this approach reduces the utility drop to 9.93% and improves UUA to 51.79%. However, it incurs a 21.04% latency overhead and increases ASR to 0.455%, as additional LLM requests introduce both computational costs and increased uncertainty. We also evaluated CSAgent’s impact on task completion steps. Overall, the impact is minimal, indicating that CSAgent’s decisions to allow or block agent actions are relatively accurate.
@@
Table 5 presents CSAgent’s per task token consumption compared to the vanilla agent. The geometric mean of consumption overhead is 35.30%, with over 96.41% consisting of prompt tokens, which are less expensive than completion tokens. Generally, more complex tasks have proportionally lower overhead. For instance, AgentBench tasks have low average token consumption, so the large context space results in higher overhead. Conversely, AndroidWorld tasks are highly complex, requiring multimodal capabilities for screen recognition, leading to very high task token consumption and lower overhead. In comparison, runtime policy generation methods introduce 1.61× additional token consumption on AgentDojo (shi2025progent). CSAgent’s low token consumption results from development-phase policy generation that applies across all devices and users without regeneration, and from requiring only one additional LLM inference per task regardless of complexity (satisfying G2).
7.5. Efficacy of Policy Evolution (RQ5)
To validate the PEF’s effectiveness, we focus on AgentDojo, where CSAgent exhibits the most significant utility degradation, and iteratively refine the corresponding context spaces. Specifically, we consolidate logs from both utility and security tests, enabling the LLM to simultaneously consider utility and security trade-offs during refinement. PEF iteratively tests and upgrades each context space until either no further improvements are identified or sufficiently high evaluation scores are achieved. In our experiments, convergence requires an average of 7.25 iterations per context space.
As shown in Table 4, CSAgent-RF maintains 0% ASR while improving UUA to 54.23%, demonstrating enhanced utility preservation without compromising security guarantees. Figure 7 shows the performance improvements: CSAgent-RF reduces utility drop to 6.76% and latency overhead to 0.53%. Notably, the reduced latency overhead primarily stems from a decrease in the average number of completion steps for some complex tasks. More precise policies enable the agent to execute tasks more efficiently by eliminating unnecessary validation failures and re-planning cycles. Our analysis reveals three primary categories of improvements in the refined context spaces. First, PEF generates more explicit and fine-grained intent definitions. Second, PEF produces more precise constraint conditions and validation logic. Third, PEF relaxes overly conservative security constraints that unnecessarily block benign operations. As a result, CSAgent-RF achieves more accurate runtime intent extraction and policy retrieval, performs more precise validation, and maintains an effective balance between blocking dangerous operations and permitting legitimate actions.
@@ 9. Related Work
Access control and contextual integrity. Traditional access control mechanisms in OS include discretionary access control (DAC) and mandatory access control (MAC) (sandhu1996authentication). SELinux  (smalley2013seandroid; wright2002selinux) exemplifies MAC principles by defining fine-grained policies that govern process interactions and resource access based on security contexts. The theory of contextual integrity (CI) (nissenbaum2004privacy; nissenbaum2009privacy), proposed in more recent years, has been explored by researchers for its applications in existing operating systems (roesner2012userdrivenac; wijesekera2015androidprem), driving the evolution of permission management and access control toward more user-transparent, secure, and flexible approaches.
@@
Access control for LLM agents. For LLM agents, access control focuses on constraining tool calls through pre-execution validation. Most approaches use contextual information to make authorization decisions, following the CI principle. AirGapAgent (bagdasarian2024airgapagent) employs a separate LLM to directly make decisions about data access permissions. GuardAgent (xiang2024guardagent) employs LLMs to generate guardrail code that validates predefined rules before executing actions. VeriSafe Agent (lee2025verisafe) describes rules through Horn clauses and verifies agent behavior based on logic-based reasoning. ShieldAgent (chen2025shieldagent) enforces explicit policy compliance by constructing action-based probabilistic rule circuits for formal verification. QuadSentinel (yang2025quadsentinelsequentsafetymachinecheckable) employs a multi-agent guard team to enforce safety through machine-checkable sequent logic rules derived from natural language policies. Conseca (tsai2025contextualagent) and Progent (shi2025progent) achieve dynamic security policy generation, enabling more autonomous responses to diverse task scenarios. AgentSentinel (hu2025agentsentinelendtoendrealtimesecurity) extends context extraction to the OS, providing real-time protection for CUAs by combining rule-based and LLM-based auditing. These approaches rely on runtime LLM-based security decision-making, suffering from security or performance limitations. AgentSpec (wang2025agentspeccustomizableruntimeenforcement) employs predefined static policies, but adopts a risk-blocking paradigm that struggles with unknown threats and lacks user intent differentiation, limiting policy flexibility.
@@
Other system-level LLM agent protection. Beyond access control, some other forms of system-level protection have been proposed for agents. The Instruction Hierarchy  (wallace2024instruction) and StruQ  (chen2024struq) implement privilege and type separation for LLM inputs to mitigate interference between different input sources. f-secure LLM system  (wu2024systemleveldefenseindirectprompt), RTBAS  (zhong2025rtbasdefendingllmagents), and SAFEFLOW  (li2025safeflowprincipledprotocoltrustworthy) introduce information flow control into LLM agents to prevent prompt injection and privacy leakage. GoEX  (patil2024goex) implements post-facto validation of agent actions and isolates dangerous operations through sandboxing. IsolateGPT  (wu2025isolategpt) employs isolation for multi-LLM application scenarios to prevent interference between LLMs and different applications. ACE (li2025acesecurityarchitecturellmintegrated) decouples agent planning into two phases to defend against untrusted third-party apps. These approaches and CSAgent can collectively form part of a defense-in-depth strategy (website:googlesecagent), working in conjunction to achieve more comprehensive protection.